OpenStudy (anonymous):

How would you work this problem? I don't understand it. Find the slope of the line whose equation is 5y=x-3 A. -3 B. -3/5 C. 1/5 D. 5

jimthompson5910 (jim_thompson5910):

divide both sides by 5 to isolate y

jimthompson5910 (jim_thompson5910):

what do you get when you do so?

OpenStudy (anonymous):

Okay I divided both sides by 5 and I got y=x-0.6

jimthompson5910 (jim_thompson5910):

\[\Large 5y=x-3 \] \[\Large y=\frac{x-3}{5} \] \[\Large y=\frac{x}{5}-\frac{3}{5} \] \[\Large y=\frac{1}{5}x-\frac{3}{5} \] So we can see the slope is 1/5

jimthompson5910 (jim_thompson5910):

you forgot to account for the x/5 portion

OpenStudy (anonymous):

Oh okay I think I understand it now thank you for the help :)